QUESTION 4

4.1 Determine L t Ht   2 ( 3)   (3)

4.2 Determine     2 3 4cosh3 t Le t . (2)

4.2 Determine 1

2

5

4 20

         

s L

s s

(5)

4.3 Determine 1

2

2 3

4 13

         

s L

s s


1
Expert's answer
2021-09-20T02:45:38-0400

4.1 "(t-3)(e^{t-3})"

"te^{t-3}-3e^{t-3}"

root (t = 3)


4.2 "\dfrac{4s-3}{s²-4s-5}"


"\dfrac{4s-3}{(s+1)(s-5)} = \dfrac7{6(s+1)} +\dfrac{17}{6(s-5)}"